OpenStudy (anonymous):

What are the dimensions of a rectangle that has a perimeter of 74 units and an area of 102 square units

OpenStudy (anonymous):

let the dimensions i.e. length and width of the rectangle be a and b respectively Thus Area of rectangle =ab i.e. 102 =ab or ab= 102--------(1 perimeter of rectangle =2(a+b) i.e. 74= 2(a+b) i.e. 2(a+b) =74 a+b= 74/2 a+b= 37--------(2 since \[(a-b)^2 =(a+b)^2 -4ab\] \[(a-b)^2 =(37)^2 -4 \times 102=1369-408=961\] \[(a-b)^2 =961 \rightarrow a-b = \sqrt{961}\] \[a-b =31---------(3\] Adding (2 and (3 we find a+b= 37 a-b =3 ------------- 2a=40 a=20 from eq (2 we find 20+b= 37 b= 37-20 b= 17 Thus the dimensions i.e. length and width of the rectangle are 20 and 17 units respectively.

OpenStudy (anonymous):

@nancycruz

OpenStudy (anonymous):

+ve solution of a2-37a+102=0